A partly full paint can has 2.75 liters of paint left in it. What is volume of the paint in cubic 

centimeters?


1
Expert's answer
2021-06-10T13:15:26-0400
"2.75\ L\cdot\dfrac{0.001\ m^3}{1\ L}\cdot\dfrac{10^6\ cm^3}{1\ m^3}=2750\ cm^3."